The sum of 2 numbers is 5. 5 times the larger number plus 4 times the smaller number is 38. Find the number
Answer by fractalier(6550)

Call the larger and smaller numbers, x and y.
Then we have
x + y = 5
and
5x + 4y = 38
A few ways to do this, so here's one...
If x + y = 5, then y = 5 - x
Now substitute and get
5x + 4(5-x) = 38
5x + 20 - 4x = 38
x + 20 = 38
x = 18
y must be 5 - 18 = -13